# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

### v1

@@ D.1 Societal Impact
Additionally, the computational efficiency provided by the adaptive event-frame fusion (FlexFuser) and frequency-adaptive learning (FAL) mechanisms reduces the need for resource-intensive training processes. This contributes to the broader societal goal of making advanced AI technologies more accessible and less energy-intensive, thereby minimizing the environmental impact of large-scale AI models. Our approach could also benefit industries beyond transportation, such as robotics for healthcare, industrial automation, and public safety.
@@ D.3 Known Limitations
While FlexEvent demonstrates significant performance improvements, there are several known limitations to our approach.
Dependence on Event Camera Quality. The effectiveness of our approach relies on the quality of the event camera sensor. Inconsistent or noisy event data, especially under poor lighting or extreme weather conditions, could affect detection performance. Future work could explore robustness to sensor noise and adaptation to diverse environmental conditions.
@@
Resource Requirements for High-Frequency Data. While FlexFuser mitigates the computational cost of training on high-frequency event data, processing extremely high-frequency event streams still requires substantial computational resources during inference. This could limit the scalability on resource-constrained devices or in real-time applications with stringent latency requirements.
Figure 7: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 8: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 9: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 10: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 11: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 12: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Appendix E E Public Resources Used

### v3

@@ D.1 Societal Impact
Additionally, the computational efficiency provided by the adaptive event-frame fusion (FlexFuse) and frequency-adaptive learning (FlexTune) mechanisms reduces the need for resource-intensive training processes. This contributes to the broader societal goal of making advanced AI technologies more accessible and less energy-intensive, thereby minimizing the environmental impact of large-scale AI models. Our approach could also benefit industries beyond transportation, such as robotics for healthcare, industrial automation, and public safety.
@@ D.3 Known Limitations
While FlexEvent demonstrates significant performance improvements, there are several known limitations to our approach, which can be summarized as follows:
Dependence on the Quality of Event Camera Data. The effectiveness of our approach relies on the quality of the event camera sensor. Inconsistent or noisy event data, especially under poor lighting or extreme weather conditions, could affect detection performance. Future work could explore robustness to sensor noise and adaptation to diverse environmental conditions.
@@
Resource Requirements for High-Frequency Data. While FlexFuse mitigates the computational cost of training on high-frequency event data, processing extremely high-frequency event streams still requires substantial computational resources during inference. This could limit the scalability on resource-constrained devices or in real-time applications with stringent latency requirements.
Appendix E Public Resources Used
@@ E.1 Public Datasets Used
  • DSEC [^2] https://dsec.ifi.uzh.ch........................................................................................................................................................................CC BY-SA 4.0
  • DSEC-Det [^3] https://github.com/uzh-rpg/dsec-det........................................................................................................................................................................GNU General Public License v3.0
  • DSEC-Detection [^4] https://github.com/abhishek1411/event-rgb-fusion........................................................................................................................................................................Creative Commons Zero v1.0 Universal
  • DSEC-MOD [^5] https://github.com/ZZY-Zhou/RENet........................................................................................................................................................................Unknown
  • Gen 1666https://www.prophesee.ai/2020/01/24/prophesee-gen1-automotive-detection-dataset........................................................................................................................................................................Prophesee Gen1 Automotive Detection Dataset License
  • 1 Mpx [^7] https://www.prophesee.ai/2020/11/24/automotive-megapixel-event-based-dataset........................................................................................................................................................................Prophesee 1Mpx Automotive Detection Dataset License
@@ E.2 Public Implementations Used
  • RVT [^8] https://github.com/uzh-rpg/RVT........................................................................................................................................................................MIT License
  • SAST [^9] https://github.com/Peterande/SAST........................................................................................................................................................................MIT License
  • SSM [^10] https://github.com/uzh-rpg/ssms_event_cameras........................................................................................................................................................................Unknown
  • LEOD [^11] https://github.com/Wuziyi616/LEOD........................................................................................................................................................................MIT License
  • DAGr [^12] https://github.com/uzh-rpg/dagr........................................................................................................................................................................GNU General Public License v3.0
  • RENet [^13] https://github.com/ZZY-Zhou/RENet........................................................................................................................................................................Unknown
Figure 9: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det. Best viewed in colors.
Figure 10: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det. Best viewed in colors.
Figure 11: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det. Best viewed in colors.
Figure 12: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 13: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 14: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Acknowledgments
The authors would like to sincerely thank the Program Chairs, Area Chairs, and Reviewers for the time and effort devoted during the review process.